OpenStudy (anonymous):

The length of a rectangle is twice the width. The area is 162yd^2. Find the length and the width. The length is ? yd. The width is ? yd.

myininaya (myininaya):

l=2w lw=162

myininaya (myininaya):

2w*w=162 2w^2=162 w^2=81 w=9 l=2(9)=18

OpenStudy (anonymous):

So the width is 9 and length is 18

OpenStudy (saifoo.khan):

Yup.

OpenStudy (anonymous):

thx